the integrating factor of the differential equation 
x (dy/dx)+2y = x2

Grade:10

3 Answers

Aditya Gupta
4 years ago
Dear student write it as dy/dx+(2/x)*y= x
Here P= 2/x so that I.F.= e^∫2/x dx
= e^2lnx
x^2
KINDLY APPROVE :))
Arun
4 years ago
Dear Meenu
 
it can be written as 
 
dy/dx + 2y/x = x
 
for integrating factor,
 
e^(integration of 2/x) dx
 
e^2 lnx
 
e^lnx2
 
x2
 
hope it helps
Vikas TU
4 years ago